### v1

# Intrinsic exceptional point – a challenge in quantum theory

Abstract
In spite of its unbroken ${\cal PT}-$symmetry, the popular imaginary cubic oscillator Hamiltonian $H^{(IC)}=p^{2}+{\rm i}x^{3}$ does not satisfy all of the necessary postulates of quantum mechanics. The failure is due to the “intrinsic exceptional point” (IEP) features of $H^{(IC)}$ and, in particular, to the phenomenon of a high-energy asymptotic parallelization of its bound-state-mimicking eigenvectors. In the paper it is argued that the operator $H^{(IC)}$ (and the like) can only be interpreted as a manifestly unphysical, singular IEP limit of a hypothetical one-parametric family of certain standard quantum Hamiltonians. For explanation, an ample use is made of perturbation theory and of multiple analogies between IEPs and conventional Kato’s exceptional points.
Keywords
quantum mechanics of unitary systems;
quasi-Hermitian Hamiltonians;
vicinity of intrinsic exceptional points;
amended perturbation theory;
imaginary cubic oscillator example;
1 Introduction
The concept of the so called “intrinsic exceptional point” (IEP) has been introduced by Siegl and Krejčiřík who, in their paper [1], studied the “prominent” imaginary cubic (IC) Schrödinger equation
$H^{(IC)}\,|\psi_{n}^{(IC)}\rangle=E_{n}^{(IC)}\,|\psi_{n}^{(IC)}\rangle\,,\ \ \ n=0,1,\ldots\,,\ \ \ H^{(IC)}=-\frac{d^{2}}{dx^{2}}+{\rm i}x^{3}\,.$ (1)
They felt motivated by the instability of the IC spectrum under perturbations [2]. They were able to complement such a numerically supported observation by several rigorous mathematical proofs (cf. also [3]). They found that “the eigenvectors of the imaginary cubic oscillator do not form a Riesz basis” [1]. In spite of having spectrum which is real, discrete and bounded below [4, 5, 6], the manifestly non-Hermitian IC Hamiltonian appeared not to be tractable, in the Riesz-basis sense, as diagonalizable.
Siegl with Krejčiřík concluded that “there is no quantum-mechanical Hamiltonian associated with it” [1]. The same authors also recalled the standard mathematical terminology as introduced in Kato’s book [7] and they reformulated their conclusion: “In the language of exceptional points, the imaginary cubic oscillator possesses an ‘intrinsic exceptional point’ ” which is, as a singularity, “much stronger than any exceptional point associated with finite Jordan blocks” [1].
These words are truly challenging, having also motivated our study of the role of IEPs in the deepest conceptual foundations of the contemporary quantum physics. It makes sense to add that Siegl with Krejčiřík only introduced the concept via the above-cited remark, i.e., without giving a formal definition. They only specified IEP as an $N=\infty$ descendant of the standard “exceptional points associated with a finite Jordan block”, i.e., of the conventional exceptional point of order $N$ (EPN). In this sense the linear differential operator $H^{(IC)}$ is really “essentially different with respect to self-adjoint Hamiltonians” [1].
The problem of interpretation of all of the non-standard, IEP-related quantum bound-state problems remains, at present, open. In our contribution to the currently running discussion of this topic (cf. also [8]) we will start, in section 2, from a brief account of what is known about the linear-algebraic EPN analogues of the ordinary differential IEP Eq. (1). We will recall the recent results of the studies of parameter-dependent $N$ by $N$ matrix quantum Hamiltonians $H^{(N)}(g)$ at finite $N<\infty$, and we will emphasize the possibility and relevance of their association with their canonical representation by an $N$ by $N$ matrix Jordan block. Thus, we will consider the toy-model quantum Hamiltonians $H^{(N)}(g)$ admitting a singular limit when $g\to g^{(EPN)}$.
In the latter limit the operator $H^{(N)}(g)$ itself ceases to be diagonalizable and, hence, it ceases to be acceptable as an eligible quantum Hamiltonian. In section 3 we will emphasize that many of its properties become really reminiscent of the IEP features of the differential-operator model (1) where the corresponding Hilbert space of states is infinite-dimensional, $N=\infty$. We will remind the readers of the existing results concerning physical meaning and impact of the EPN-related finite-dimensional models. We will explain that in many (often called “quasi-Hermitian” [9, 10]) quantum models of such a type the limiting transition $g\to g^{(EPN)}$ can be interpreted as one of the most natural realizations of a genuine quantum phase transition (cf., e.g., the description of a class of exactly solvable models of such a process in [11]).
In section 4 the emphasis will be shifted to the $N\to\infty$ scenarios and to the existence of several very useful analogies between both of the IEP and EPN singular extremes. We will point out that in such a comparison the key role of a methodical guide may be expected to be played by (possibly, amended) perturbation theory. Interested readers will be recommended to find a phenomenologically oriented inspiration as well as many related technical details in the older paper [12]. What is important is that the authors studied there a fairly realistic non-Hermitian Hamiltonian describing an N-particle Bose-Einstein condensate generated by a sink and a source in interaction. Using a combination of several complementary numerical as well as analytic and perturbation methods they managed to detect the presence of the EPN singularities in their model. They also revealed and explained that under small perturbations these singularities did unfold in a very specific manner. Their results appeared encouraging because, as they mentioned, the EPN-related problems are truly challenging and their “further investigations …remain tasks for future research.” In our present paper, we just decided to follow the recommendation.
In sections number 5 and 6 we will, in particular, address our main technical task, viz., an IEP-related generalization of the well known perturbation-theory-based description of a generic unitary quantum system near its singularity. We will succeed in showing that many known technical tricks used and tested near EPN at $N<\infty$ can immediately be transferred to the technically more difficult quantum-dynamical scenarios in which the generic Hamiltonians $H(g)$ lie very close to an IEP limiting extreme.
Ultimately, in our last two sections 7 and 8 and also in the series of six brief Appendices we will complement our considerations by several quantum-physics-oriented contextual remarks.
2 Conventional exceptional points associated with finite Jordan blocks
In [1] we read that the existence of the IEP singularity “does not restrict to the particular Hamiltonian” of Eq. (1) so that some “new directions in physical interpretation” of all of the analogous non-Hermitian quantum models have to be sought “since their properties are essentially different with respect to self-adjoint Hamiltonians” [1].
This makes the IC model important as a genuine methodical as well as conceptual challenge. Here, we intend to propose and advocate the idea that the resolution of the problem could be guided by the EPN-related “good basis” problem and by the existence of parallels between quantum systems near their respective IEP and EPN singularities.
The study of these parallels could proceed in several independent directions (cf. the samples of some of them in [8] or in [13]). In what follows, we will explain that and how one of these directions could make use of perturbation-expansion techniques.
2.1 The phenomenon of EPN degeneracy
In review paper [14] the very first word of Abstract emphasizes that every operator $H$ eligible as an observable Hamiltonian of a unitary quantum system in Schrödinger picture [15] must be diagonalizable. For any specific one-parametric family of Hamiltonians $H(g)$ such a requirement is not satisfied in the EPN limit $g\to g^{(EPN)}$. Then, the operator can consistently be treated as Hamiltonian only when $g\neq g^{(EPN)}$.
In an opposite direction of argumentation one could recall the existence of exactly solvable quasi-Hermitian $N$ by $N$ matrix models $H^{(N)}(g)$ of paper [11] for which there exists a vicinity of $g^{(EPN)}$ (say, ${\cal D}$) inside which the respective quantum system is found to admit the standard physical probabilistic interpretation. The diagonalizability of the Hamiltonians $H^{(N)}(g)$ plus a few other technical assumptions (concerning the spectrum) then imply that, in principle at least, we may construct all of the bound-state solutions of the so called time-independent Schrödinger equation
$H^{(N)}({g})\,|\psi_{n}(g)\rangle=|\psi_{n}(g)\rangle\,{E}^{(N)}_{n}(g)\,,\ \ \ \ \ n=0,1,\ldots,N-1\,.$ (2)
Now, whenever the dimension of the Hilbert space of states is finite, $N<\infty$, we may immediately notice that even in the generic non-degenerate case, all of the eigenvalues ${E}^{(N)}_{n}(g)$ and eigenvectors $|\psi_{n}(g)\rangle$ with $g\in{\cal D}$ degenerate in the ultimate (albeit manifestly unphysical) EPN limit,
$\lim_{g\to g^{(EPN)}}\,{E}^{(N)}_{n}(g)=E^{(EPN)}\,,\ \ \ \ \lim_{g\to g^{(EPN)}}\,|\psi_{n}(g)\rangle=|\Psi^{(EPN)}\rangle\,,\ \ \ \ n=0,1,\ldots,N-1\,.$ (3)
This leads to the following observations:
  • [1] for all of the “acceptable” $g\neq g^{(EPN)}$ lying in a vicinity of the EPN value, $g\in{\cal D}$, the normalized eigenvectors $|\psi_{n}(g)\rangle$ of $H^{(N)}(g)$ are getting almost parallel to each other.
  • [2] at the “unacceptable” value of $g=g^{(EPN)}\notin{\cal D}$ their set ceases to serve as a basis suitable, say, for the purposes of perturbation theory.
  • [3] at $g=g^{(EPN)}$ one can still construct a “good basis” composed of the single remaining (degenerate) eigenvector $|\psi_{0}(g^{(EPN)})\rangle=|\Psi_{0}\rangle$ and of an $(N-1)-$plet of linearly independent associated vectors $|\Psi_{j}\rangle$ with $j=1,2,\ldots,N-1$.
2.2 EPN and modified Schrödinger equation
The latter “good basis” can be perceived as an $N-$plet of column vectors. They may be arranged into the following formal $N$ by $N$ matrix,
$\{|\Psi_{0}\rangle,|\Psi_{1}\rangle,\ldots,|\Psi_{N-1}\rangle\}:=R^{(EPN)}$ (4)
called, usually, transition matrix. Thus, we may introduce the two-diagonal Jordan block
$J^{(N)}\left(\eta\right)=\left(\begin{array}[]{ccccc}\eta&1&0&\ldots&0\\ 0&\eta&1&\ddots&\vdots\\ 0&0&\eta&\ddots&0\\ \vdots&\ddots&\ddots&\ddots&1\\ 0&\ldots&0&0&\eta\end{array}\right)\,$ (5)
and define the transition matrix as solution of the following Schrödinger-like equation
$H^{(N)}({g}^{(EPN)})\,{R}^{(EPN)}={R}^{(EPN)}\,{J}^{(N)}(E^{(EPN)})\,.$ (6)
Interested readers are recommended to find a constructive illustration of the reconstruction of transition matrix ${R}^{(EPN)}$ from the Hamiltonian in [16] where the illustrative solvable Hamiltonians were real matrices which were tridiagonal and multiparametric: At $N=2J$ one had
$H^{(2{J})}(a,b,\ldots,z)=\left[\begin{array}[]{cccccccc}2{J}-1&z&0&\ldots&&&&\\ -z&\ddots&\ddots&\ddots&\vdots&&&\\ 0&\ddots&3&b&0&\ldots&&\\ \vdots&\ddots&-b&1&a&0&\ldots&\\ &\ldots&0&-a&-1&b&0&\ldots\\ &&\ldots&0&-b&-3&\ddots&\\ &&&\vdots&\ddots&\ddots&\ddots&z\\ &&&&\ldots&0&-z&1-2{J}\end{array}\right]$ (7)
etc (for a few further related comments see also Appendices A.1 and A.2 below).
3 The mechanism of unfolding of the EPN degeneracy
3.1 Hypothesis of admissibility of the choice of $g\neq{g}^{(EPN)}$
The motivation of the above-outlined amendment of the basis is twofold. Firstly, it enables us to re-read our Schrödinger-like Eq. (6) as an equivalent linear-algebraic relation
$\left[{R}^{(EPN)}\right]^{-1}\,H({g}^{(EPN)})\,{R}^{(EPN)}={J}^{(N)}({E}^{(EPN)})\,$ (8)
i.e., as a definition of a canonical $N$ by $N$ matrix Jordan-block representation of any non-Hermitian Hamiltonian of interest at its EPN singularity. Secondly, the amended basis might find an application in a reformulation of standard perturbation theory. In such a reformulation, the role of the unperturbed Hamiltonian could have been played by its unphysical, singular EPN limit. The trick is that we use the columns of ${R}^{(EPN)}$ as unperturbed basis, assuming only that, in the overall perturbation-theory spirit, the perturbed system acquires a standard phenomenological interpretation for the parameters $g$ lying inside a suitable “physical” vicinity ${\cal D}$ of the EPN singularity.
The latter philosophy is to be advocated and used in what follows. We will only assume the knowledge of the transition matrix ${R}^{(EPN)}$ at an exceptional point of finite order and we will only extend the use of this basis also to some vicinity of the singularity. This will enable us to invert the limiting process $g\to{g}^{(EPN)}$ and to consider the original Hamiltonians at some $g\neq{g}^{(EPN)}$. Our knowledge of transition matrix will yield the model in a simplified form described as a perturbation of a Jordan block matrix,
$\left[{R}^{(EPN)}\right]^{-1}\,H^{(N)}({g})\,{R}^{(EPN)}={J}^{(N)}({E}^{(EPN)})+\lambda\,V^{(N)}({g})\,.$ (9)
A priori, we will only have to demand that the auxiliary variable $\lambda={\cal O}({g}-{g}^{(EPN)})$ remains small.
3.2 The possibility of keeping the perturbed spectrum real
In papers [17, 18] we considered the above-mentioned quantum-dynamics scenarios and, thoroughly, we studied there the criteria of smallness of the perturbations $V^{(N)}$. We showed that the conditions of the stability and unitarity of the system can be given a mathematically as well as phenomenologically consistent form.
For an illustration of details let us set, conveniently, $E^{(EPN)}=0$ and let us consider the bound-state problem as a perturbation of its EPN limit,
$\left[{J}^{(N)}(0)+\lambda\,V^{(N)}\right]\,|{\Psi}(\lambda)\rangle=\epsilon(\lambda)\,|{\Psi}(\lambda)\rangle\,$ (10)
with the energy levels counted, whenever needed, by a subscript or superscript. We will never use this index, nevertheless, keeping it just dummy. We will rather introduce another subscript which will run, say, from $1$ to $N$ and which will number the components $\Psi_{j}$ of the ket vector $|{\Psi}\rangle$ (here we are also dropping the argument $\lambda$ as redundant).
This convention enables us to fix the norm of $|{\Psi}\rangle$ by the choice of ${\Psi}_{1}=1$ and to define another, “shifted” column vector
$\left(\begin{array}[]{c}{\Psi}_{2}\\ {\Psi}_{3}\\ \vdots\\ {\Psi}_{N}\\ \Omega_{N}\end{array}\right):=\left(\begin{array}[]{c}y_{1}\\ y_{2}\\ \vdots\\ y_{N-1}\\ y_{N}\end{array}\right)=\vec{y}\,$ (11)
where $\Omega_{N}$ is a new auxiliary variable. Next we notice that the $N$ by $N$ matrix
$A=A(N,\epsilon)=\left(\begin{array}[]{ccccc}1&0&0&\ldots&0\\ \epsilon&1&0&\ddots&\vdots\\ \epsilon^{2}&\epsilon&\ddots&\ddots&0\\ \vdots&\ddots&\ddots&1&0\\ \epsilon^{N-1}&\ldots&\epsilon^{2}&\epsilon&1\end{array}\right)\,$ (12)
is just an inverse of the two-diagonal matrix
$A^{-1}=\left(\begin{array}[]{ccccc}1&0&0&\ldots&0\\ -\epsilon&1&0&\ddots&\vdots\\ 0&-\epsilon&\ddots&\ddots&0\\ \vdots&\ddots&\ddots&1&0\\ 0&\ldots&0&-\epsilon&1\end{array}\right)\,.$ (13)
Finally we select the first column of the matrix in Eq. (10) and denote it by another dedicated symbol,
$\left(\begin{array}[]{c}\epsilon-\lambda\,{V}_{1,1}\\ -\lambda\,{V}_{2,1}\\ \vdots\\ -\lambda\,{V}_{N,1}\end{array}\right):=\vec{r}=\vec{r}(\lambda)\,.$ (14)
All of these abbreviations convert our initial homogeneous Schrödinger Eq. (10) into its equivalent matrix form
$(A^{-1}+\lambda\,Z)\,\vec{y}=\vec{r}\,$ (15)
or, better,
$(I+\lambda\,A\,Z)\,\vec{y}=A\,\vec{r}\,$ (16)
where the symbol $Z$ stands just for a modified form of the matrix of perturbation,
$V^{(N)}\ \to\ Z=\left(\begin{array}[]{ccccc}{V}_{1,2}&{V}_{1,3}&\ldots&{V}_{1,N}&0\\ {V}_{2,2}&{V}_{2,3}&\ldots&{V}_{2,N}&0\\ \ldots&\ldots&\ldots&\vdots&\vdots\\ {V}_{N,2}&{V}_{N,3}&\ldots&{V}_{N,N}&0\end{array}\right)\,.$ (17)
In paper [17] we proved that the construction of the perturbation corrections now becomes reduced to the self-consistency condition
$\Omega_{N}=0\,.$ (18)
In the same reference, interested readers may also find an explicit form of the construction in the leading-order approximation. In fact, its basic aspects are also worth recalling because they immediately help to clarify the meaning of the rather vague above-mentioned assumption of the “sufficient smallness” of the perturbation. In essence, it is sufficient to employ the Taylor-series expansion of the resolvent which yields formula
$\vec{y}^{(solution)}(\epsilon)=A\,\vec{r}-\lambda\,A\,\,Z\,A\,\vec{r}+\lambda^{2}\,A\,\,Z\,A\,Z\,A\,\vec{r}-\ldots\,.$ (19)
Such a form of the wave-function-representing ket-vector still depends on the variable parameter $\epsilon$ but, ultimately, all of the eligible values of $\epsilon$ become fixed by constraint (18).
The latter constraint plays the role of secular equation which has, after all, just the scalar vector-component form
${y_{N}}^{(solution)}(\epsilon)=0\,.$ (20)
In the last step of the construction, obviously, we just have to solve such an explicit transcendental equation in order to get all of the alternative perturbation-generated energy corrections $\epsilon$. In a direct dependence on the model in question, precisely the study of the roots of this equation also offers the criterion of the reality of the whole spectrum in the leading-order approximation.
4 Limit $N\to\infty$ and the anomalous IEP Hamiltonians
The message to be extracted from the above-outlined EPN-based construction is that for the purposes of transition to its IEP analogue we may try to make use of the IEP - EPN similarities. The main one will consist in the unperturbed-Hamiltonian interpretation of the singular IEP operator tractable, in some sense, as a large-$N$ descendant of its finite-$N$ EPN analogues.
4.1 The IEP-related asymptotic degeneracy of eigenvectors
In the analysis of IEP singularities a central role is certainly played by the phenomenon of the asymptotic confluence of infinitely many eigenvectors, not accompanied by the confluence of the eigenvalues. Indeed, in [8] we read that “for matrices approaching an exceptional point it is known [19] that the corresponding eigenvectors are tending to coalesce. For the infinite-dimensional Hilbert space (and Krein space) setup of the IC model, the eigenfunctions of the Hamiltonian having diverging projector norms and asymptotically approaching a PT phase transition region at spectral infinity signal a possible tendency toward collinearity and isotropy of an infinite number of these eigenfunctions” [8].
The EPN - IEP parallels are certainly incomplete so that in the latter case, an amendment of the notation might prove useful. The first mathematical subtlety which we will have to keep in mind is that for a generic IEP model the spectrum itself remains non-degenerate so that, in a way sampled by the IC example, the generic IEP Schrödinger equation
$H^{(IEP)}\,|\psi_{n}^{(IEP)}\rangle=E_{n}^{(IEP)}\,|\psi_{n}^{(IEP)}\rangle\,,\ \ \ n=0,1,\ldots\,$ (21)
can still be considered analogous to its EPN partners. Once the spectrum is found real and discrete (which is precisely the case of the illustrative IC Schrödinger Eq. (1)), the same property also characterizes the formally independent Hermitian conjugate Schrödinger equation problem
$\left[H^{(IEP)}\right]^{\dagger}\,|\psi_{n}^{(IEP)}\rangle\!\rangle=E_{n}^{(IEP)}\,|\psi_{n}^{(IEP)}\rangle\!\rangle\,,\ \ \ n=0,1,\ldots\,.$ (22)
The IEP-characterizing phenomenon of asymptotic degeneracy enables us to re-establish the above-mentioned analogy with the EPN form of confluence of the eigenfunctions. This phenomenon involves, first of all, the right eigenvectors $|\psi_{n}^{(IEP)}\rangle$ of $H^{(IEP)}$. For them we have
$|\psi_{M+k}^{(IEP)}\rangle\approx|\psi_{M+k+1}^{(IEP)}\rangle\,,\ \ \ \ k=1,2,\ldots\,$ (23)
at $M\gg 1$. Similarly, the degeneracy concerns also the left alias  “brabra” [20] eigenvectors $\langle\!\langle\psi_{n}^{(IEP)}|$ of the same $H^{(IEP)}$. Often we rather refer to their conjugate form $|\psi_{n}^{(IEP)}\rangle\!\rangle$ of “ketket” eigenvectors of conjugate $\left[H^{(IEP)}\right]^{\dagger}$. In this representation we have
$|\psi_{M+k}^{(IEP)}\rangle\!\rangle\approx|\psi_{M+k+1}^{(IEP)}\rangle\!\rangle\,,\ \ \ \ k=1,2,\ldots\,.$ (24)
In both Eqs. (23) and (24) the degree of confluence depends on the Hamiltonian and it may be expected to grow with the growth of $M$.
The phenomenon of the confluences (23) and (24) finds its formal predecessor in the finite-dimensional case in which, during the transition to singularity $g\to g^{(EPN)}$ the elements of the $N-$plet of eigenvectors of any preselected $N$ by $N$ Hamiltonian matrix $H=H^{(N)}(g)$ really lose their mutual linear independence. Still, the analogy of a genuine IEP system with the IEP-mimicking $N=\infty$ EPN extreme is incomplete since in the former case the spectrum remains non-degenerate. A more explicit analysis is necessary.
4.2 Canonical representation of $H^{(IEP)}$
The IEP-characterizing non-degeneracy of eigenvalues can be perceived as a serendipitious simplification of their study. Still, a decisive IEP-related difficulty results from the effective asymptotic parallelization of the unlimited number of eigenvectors.
This forces us to recall, as our main source of inspiration, relations (5) and (6) of section 2 above. In the IEP case our key task can be now identified as an appropriate generalization of the transition matrices ${R}^{(EPN)}$ which played key role in the perturbation-theory considerations of section 3. In other words, we have to replace Eq. (6) by a modified eigenvalue-like problem
$H^{(IEP)}\,{\cal R}^{(IEP)}={\cal R}^{(IEP)}\,{\cal J}^{(IEP)}\,$ (25)
in which the low-lying eigenstates do not require any specific attention. Thus, the conventional Jordan-block-like bidiagonal (i.e., minimally non-diagonal) canonical-matrix structure of Eq. (5) will only reemerge here in a infinite-dimensional submatrix of upgraded
${\cal J}^{(IEP)}=\left(\begin{array}[]{cccc|cccc}E_{0}&0&\ldots&0&0&\ldots&&\\ 0&E_{1}&\ddots&\vdots&\vdots&&&\\ \vdots&\ddots&\ddots&0&0&\ldots&&\\ 0&\ldots&0&E_{K-1}&0&0&\ldots&\\ \hline\cr 0&\ldots&0&0&E_{K}&1&0&\ldots\\ \vdots&&\vdots&0&0&E_{K+1}&1&\ddots\\ &&&\vdots&0&0&E_{K+2}&\ddots\\ &&&&\vdots&\vdots&\ddots&\ddots\\ \end{array}\right)\,.$ (26)
In this arrangement the partitioning of the basis may be characterized by the respective projectors $P$ (on the first $K$ lowest eigenstates) and $Q=I-P$. In a certain parallel with EPN, a key technical challenge will now lie in a weakening or removal of the asymptotic degeneracies (23) and (24).
5 Towards a regularization of $H^{(IEP)}$s by perturbation
From a purely historical point of view the idea of “prominence” of the IEP-sampling Schrödinger Eq. (1) dates back to its methodical role in field theory [21] and to the Bessis’ and Zinn-Justin’s empirically revealed conjecture (cf. [4], cited also in [5]) that the spectrum $\{E_{n}^{(IC)}\}$ of $H^{(IC)}$ is real, discrete and bounded from below, i.e., tractable, in principle at least, as the set of the observable energy levels of bound states of a unitary quantum system. In spite of the manifest non-Hermiticity of Hamiltonian $H^{(IC)}$, the model was temporarily accepted as potentially compatible with all of the principles and postulates of quantum mechanics. The corresponding technical details may be found in review paper [22].
Unfortunately, the end of the excitement came after the Siegl’s and Krejčiřík’s rigorous proof that the IC model cannot in fact be assigned any form of conventional probabilistic interpretation in a mathematically consistent manner [1]. More or less the same conclusion has been also made, in a not yet published preprint, by Günther and Stefani [8]. At present, in the context of the unitary-evolution part of non-Hermitian quantum mechanics the problem of a correct physical interpretation of the IC model (1) remains, therefore, unresolved.
Concerning the future developments, we remain a bit skeptical because the IEP nature of the IC model looks, in many a respect, only too similar to its much better understood (and manifestly singular and unphysical) EPN-related finite-dimensional analogues.
5.1 The IEP - EPN differences and parallels
The essence of the anomalous nature of any IEP-related Hamiltonian $H^{(IEP)}$ lies in the asymptotic degeneracies (23) and (24) of its respective right and left eigenvectors. At the same time, a weak point of the amendment of the basis as mediated by the choice of non-diagonal matrix (26) may be seen in the necessity of specification of a “sufficiently large” onset $K\gg 1$ of the de-parallelization. Such a specification is just numerically, computer-precision motivated. In contrast to the above-outlined treatment of the EPN scenarios where the dimension $N$ was fixed, the IEP-implied choice of any finite $K$ is purely pragmatic and virtually arbitrary.
We now intend to show that, surprisingly enough, the apparently more or less accidental flexibility of our choice of $K$ can in fact become an important tool facilitating an EPN-resembling regularization and consistent interpretation of quantum systems near their IEP singular extreme.
First of all, there is no doubt about the necessity of transition from the less suitable unperturbed basis (composed of eigenvectors) to an “anomalous” basis resembling Eq. (4). The reason is the situation as represented by Eqs. (25) and (26): only a rectification of the underlying biorthogonal or biorthonormal basis [23] can re-establish the EPN - IEP parallels even when achieved, also in the latter case, at an expense of the non-diagonality and non-Hermiticity of matrix (26).
Although the EPN singularity encountered at finite matrix dimensions $N<\infty$ is, according to paper [1], perceivably weaker than its IEP analogue, the essence of our present message will be complementary. Basically, we will emphasize that one can also reveal and make a productive use of certain partial similarities between the two concepts. In particular, we propose that the above-outlined possibility and feasibility of treating the manifestly unphysical finite-dimensional singular matrices $H^{(N)}(g^{(EPN)})$ as formally acceptable unperturbed Hamiltonians is to be transferred also to the IEP context. Moreover, also the construction of the equally anomalous EPN-related “good” basis (4) composed of the columns of the transition matrix should be, mutatis mutandis, introduced and reconstructed from any given anomalous Hamiltonian $H^{(IEP)}$.
5.2 IEP-unfolding bases
The IEP (i.e., $N=\infty$) and EPN (i.e., $N<\infty$) singularities share the phenomenon of the parallelization of eigenvectors. In a small vicinity of the singularity the analysis has to rely upon a properly adapted form of perturbation-theory. Our present proposal of transfer of this idea from EPN to IEP will be inspired, therefore, by section 3.
The parallels are, naturally, incomplete so that in the IEP setting certain truly specific features have to be expected to emerge. For the purposes of clarification let us mention that even if we fix a finite $K\gg 1$ it remains far from obvious how to follow the analogy with Eqs. (2) and (4) and how to treat also $H^{(IEP)}$ as an unperturbed Hamiltonian. The reason is that we do not have any immediate analogue of Eq. (9). In the models as sampled by the IC oscillator we also miss a parameter $g$ or $\lambda$ which would control the form and size of perturbations needed for a phenomenologically motivated unfolding of the manifestly unphysical IEP singularity.
This being admitted, we may still be guided by the EPN dynamical scenario as outlined in the preceding sections 2 and 3. In the study of the IEP systems, first of all, we should construct a good unperturbed basis in Hilbert space, therefore. The most natural IEP analogue of the EPN-related Jordan-block-matrix (5) is to be seen in its IEP-related amendment (26), rendering the EPN-related unperturbed Schrödinger-like Eq. (6) replaced by its IEP-related alternative (25).
In connection with the standard and unmodified conjugate eigenvalue problems (21) and (22) the difficulty is that in the Siegl’s and Krejčiřík’s words “the eigenvectors, despite possibly being complete, do not form a ‘good’ basis, i.e., an unconditional (Riesz) basis” [1]. Thus, the left and right eigenvectors of $H^{(IEP)}$ can only be used as a basis in the $P-$projected subspace of the Hilbert space. Otherwise, the EPN - IEP parallelism has to be fully taken into account, i.e., in Eq. (25), one has to recall the EPN-related definition (4) and define, in Eq. (25), its present IEP-related calligraphic-symbol partner ${\cal R}^{(IEP)}$ as the following concatenated infinite set of column vectors
${\cal R}^{(IEP)}=\{|\psi_{0}\rangle,|\psi_{1}\rangle,\ldots,|\psi_{K-1}\rangle,|f_{K}\rangle,|f_{K+1}\rangle,|f_{K+2}\rangle,\ldots\}$ (27)
composed of the mere first $K$ eigenkets $|\psi_{j}\rangle$ complemented by the modified, associated-like ket vectors $|f_{K+k}\rangle$ with $k=0,1,2,\ldots$.
5.3 Recurrences
The possibility (and also, in some sense, the necessity) of the explicit construction of the latter subfamily of the new ket vectors is in fact the very core of our present innovation of the foundations of the formalism of quantum mechanics. Briefly, our basic message is that in a way which parallels the EPN-related considerations of section 3 above, our present introduction of the nontrivial IEP-motivated transition matrix (27) may be expected to play a key role in the regularization of any singular $H^{IEP)}$ via its suitable small perturbations.
The replacement of eigenvectors $|\psi_{K+k}\rangle$ by non-eigenvectors $|f_{K+k}\rangle$ in (27) has to weaken the asymptotically increasing parallelism between the subsequent columns of the transition matrix ${\cal R}^{(IEP)}$. The infinite-dimensional matrix form of transition matrix (27) makes this task different from its EPN predecessor. In technical terms, the insertion of (27) may be used to reduce the nontrivial part of Eq. (25) to the sequence of recurrences
$\left(H^{(IEP)}-E^{(IEP)}_{K+m}\right)\,|f_{K+m}\rangle=|f_{K+m-1}\rangle\,,\ \ \ \ m=1,2,\ldots\,$ (28)
with the initial choice of $|f_{K}\rangle=c_{0,0}|\psi_{K}\rangle$ using any $c_{0,0}\neq 0$.
The solution of these relations can be then given the form of finite sum
$|f_{K+p}\rangle=\sum_{n=0}^{p}\,c_{p,n}\,|\psi_{K+n}\rangle\,,\ \ \ \ p=0,1,\ldots$ (29)
where the leading coefficient $c_{k,k}$ is arbitrary. Now we assume and recall the biorthonormality of the eigenbasis yielding $\langle\!\langle\psi_{m}|\psi_{n}\rangle=\delta_{m,n}$ and enabling us to convert Eq. (28), i.e., the recurrences for kets into the recurrences for coefficients,
$c_{k,m}=(E_{K+m}-E_{K+k})^{-1}c_{k-1,m}\,,\ \ \ \ \ m=0,1,\ldots,k-1\,,\ \ \ \ \ k=1,2,\ldots\,.$ (30)
Our freedom of choice of the highest-component coefficients $c_{k,k}$ enables us to suppress the IEP-accompanying asymptotic parallelization of the vectors of basis in Hilbert space. The goal is achieved. For every particular IEP model we recall recurrences (30) and replace the $Q-$projected part of the basis composed of eigenvectors by the $Q-$projected part of the basis composed, up to the first item $|f_{K}\rangle=c_{0,0}|\psi_{K}\rangle$, of non-eigenvectors. And this is precisely what has been done in Eq. (27).
6 Constructive IEP-perturbation considerations
6.1 Formulation of the problem
Günther with Stefani [8] stated that “what is still lacking” in the IEP setup “is a simple physical explanation scheme for the non-Rieszian behavior of the eigenfunction sets”. We agree. We are even more skeptical because we would rather say that the expected ‘simple physical explanation’ making, in particular, the popular IC oscillators (1) physical need not exist at all. Indeed, we believe that a consistent physical closed quantum system interpretation could much more easily be assigned to suitable perturbations of the “seed” IEP oscillators with uncertain interpretation (cf. [24] in this respect).
Our belief is supported by the existence of parallels between the IEP and EPN scenarios. On this ground one could really become able to assign a sound phenomenological meaning to many hypothetical parameter-dependent Hamiltonians ${\cal H}^{(new)}({\lambda})$ defined as certain “admissible” perturbations of the extreme IEP reference operators
$H^{(IEP)}\,\equiv\,{\cal H}^{(new)}({0})$
(see, in this respect, also the methodical guidance as provided by the illustrative EPN-related Eq. (42) in section 7 below).
Open questions emerge when we fix a sufficiently large $K$, separate the Hilbert space of states into its two more or less decoupled subspaces and when we finally introduce a hypothetical perturbed Hamiltonian ${\cal H}^{(new)}(\lambda)$ and the following IEP analogue of Eq. (9),
$\left[{\cal R}^{(IEP)}\right]^{-1}\,{\cal H}^{(new)}({\lambda})\,{\cal R}^{(IEP)}={\cal J}^{(IEP)}+\lambda\,{\cal V}\,.$ (31)
The analogy with EPNs is incomplete because here, the spectrum of the unperturbed zero-order Hamiltonian remains non-degenerate. This is a simplification which can be perceived as partially compensating the increase of the overall complexity of the IEP problem.
Incidentally, a similar simplification has also been detected in the realistic EPN-supporting Bose-Hubbard model of paper [12] where, in dependence on parameters, the authors had to use both  the degenerate and non-degenerate versions of perturbation theory. Thus, no abstract conceptual problems have to be expected to emerge after one returns to the generic IEP-related dynamics. Still, as long as the IEP-related problems are infinite-dimensional, the perturbed IEP spectrum cannot be deduced from any analogue of the EPN-based implicit-definition constraint (18). The study of properties of the vicinity of the IEP singularity cannot be based on a direct reference to its EPN analogue.
6.2 Structure of solutions
In the light of the $P+Q$ partitioning of matrix ${\cal J}^{(IEP)}$ of Eq. (26) the constructiton of the perturbed forms of the low-lying bound states remains standard. The $P-$projected states may be ignored just as certain decoupled observers. Only the treatment of the “asymptotic”, $Q-$projected components of the quantum system in question becomes difficult and singular, “for instance, due to spectral instabilities” [1].
This leads to the necessity of solving the perturbed Schrödinger equation
$\left[{\cal J}^{(IEP)}+\lambda\,{\cal V}\right]\,|\psi(\lambda)\rangle=E(\lambda)\,|\psi(\lambda)\rangle$ (32)
where $\lambda\neq 0$ (so that we avoid the IEP singularity) and where we have to insert
$|\psi(\lambda)\rangle=|\psi(0)\rangle+\lambda\,|\psi^{[1]}\rangle+\lambda^{2}\,|\psi^{[2]}\rangle+\ldots$ (33)
and
$E(\lambda)=E(0)+\lambda\,E^{[1]}+\lambda^{2}\,E^{[2]}\rangle+\ldots\,.$ (34)
In the light of definition (26) the $P-$projected part of our unperturbed Hamiltonian ${\cal J}^{(IEP)}$ is a diagonal matrix containing the unperturbed bound state energy eigenvalues $E_{0}$, $E_{1}$, …$E_{K}$. All of the related perturbed low-lying bound states can be then constructed using the conventional Rayleigh-Schrödinger perturbation theory of textbooks [15]. For any practical purposes it is fully acceptable just to make the choice of a sufficiently large dimension $K$ of the $P-$projected subspace, therefore.
In our present, conceptually more ambitious analysis of the problem it makes sense to turn attention to the states with the high-lying zero-order unperturbed energies. The related unperturbed ket vectors $|\psi(0)\rangle$ will lie in the complementary (and infinite-dimensional) $Q-$projected subspace of Hilbert space. The more or less conventional construction of its perturbed descendant of Eq. (33) will then possess several anomalous features of course.
The first anomaly is that the $Q-$projected part $Q{\cal J}^{(IEP)}Q$ of our unperturbed Hamiltonian is manifestly non-Hermitian. Even after a tentative finite-matrix truncation of the perturbed eigenvalue problem using a sufficiently large cut-off $N\gg K\gg 1$ of the Hilbert space bases, the implementation of the conventional Rayleigh-Schrödinger recipe would require a spectral representation of the unperturbed Hamiltonian operator.
Due to the specific upper-triangular two-diagonal structure of matrix $Q{\cal J}^{(IEP)}Q$, the construction of a biorthonormalized basis would be needed. Thus, the left eigenvectors of $Q{\cal J}^{(IEP)}Q$ (i.e., in the notation of paper [20], ketkets, $|\chi_{j}\rangle\!\rangle$) will be complicated and different from their right-eigenvector biorthogonal partners $|\chi_{j}\rangle$. This means that also the conventional Rayleigh-Schrödinger elementary unperturbed projectors $|\chi_{j}\rangle\!\rangle\langle\!\langle\chi_{j}|$ (needed during the construction) will have a practically prohibitively complicated explicit matrix structure.
One could also find another, more immediate indication of the possible emergence of irregularities in section 3 where Eq. (20) playing the role of an ultimate transcendental equation determining all of the perturbed EPN eigenvalues was just a constraint imposed upon the very last, $N-$th component of a relevant ket vector. Needless to add that in the IEP setting one should consider $N\to\infty$ so that the direct analogy with EPNs gets broken.
Another, independent word of warning might originate from the fact that for all of the truly high energy levels $E_{K+m}(\lambda)$ with $m\gg 1$ the use of the explicit Rayleigh-Schrödinger recipe would require the derivation of formulae which would be $m-$dependent and different for different $m-$numbered excitations. Fortunately, the latter technical obstacle and difficulty has a comparatively elementary resolution because what is fully at our disposal is our choice of the value of $K$. We may feel free to work, exclusively, with the properly innovated Rayleigh-Schrödinger formulae deduced just at a single value of $m$, i.e., say, at $m=0$.
This certainly simplifies our task. In methodical setting, it will be sufficient to work with the Hamiltonian of Eq. (26) at $K=0$. Thus, one just has to solve Schrödinger equation
$\left[\left(\begin{array}[]{cccc}E_{0}-E(\lambda)&1&0&\ldots\\ 0&E_{1}-E(\lambda)&1&\ddots\\ 0&0&E_{2}-E(\lambda)&\ddots\\ \vdots&\ddots&\ddots&\ddots\\ \end{array}\right)+\lambda\,\left(\begin{array}[]{cccc}{\cal V}_{00}&{\cal V}_{01}&{\cal V}_{02}&\ldots\\ {\cal V}_{10}&{\cal V}_{11}&{\cal V}_{12}&\ddots\\ {\cal V}_{201}&{\cal V}_{21}&{\cal V}_{22}&\ddots\\ \vdots&\ddots&\ddots&\ddots\\ \end{array}\right)\right]\,|\psi(\lambda)\rangle=0\,$ (35)
where
$|\psi(\lambda)\rangle=\left(\begin{array}[]{c}\psi_{0}(0)\\ \psi_{1}(0)\\ \vdots\end{array}\right)+\sum_{k=1}^{\infty}\lambda^{k}\,\left(\begin{array}[]{c}\psi_{0}^{[k]}\\ \psi_{1}^{[k]}\\ \vdots\end{array}\right)\,.$ (36)
As long as $E_{0}=E(0)$ and $|\psi_{j}(0)\rangle=0$ at all $j\neq 0$, it makes sense to abbreviate $E_{k}-E(\lambda):=F_{k}(\lambda)$ and remember that $F_{0}(\lambda)=\lambda E^{[1]}+{\cal O}(\lambda^{2})$.
In the first-order approximation we have, therefore, equation
$\left(\begin{array}[]{cccc}-\lambda E^{[1]}&1&0&\ldots\\ 0&F_{1}(0)&1&\ddots\\ 0&0&F_{2}(0)&\ddots\\ \vdots&\ddots&\ddots&\ddots\\ \end{array}\right)\left[\left(\begin{array}[]{c}\psi_{0}(0)\\ 0\\ 0\\ \vdots\end{array}\right)+\lambda\,\left(\begin{array}[]{c}\psi_{0}^{[1]}\\ \psi_{1}^{[1]}\\ \psi_{2}^{[1]}\\ \vdots\end{array}\right)\right]=$ (37)
$=\lambda\,\left(\begin{array}[]{c}-E^{[1]}\,\psi_{0}(0)\\ 0\\ 0\\ \vdots\end{array}\right)+\lambda\,\left(\begin{array}[]{cccc}1&0&o&\ldots\\ F_{1}(0)&1&0&\ddots\\ 0&F_{2}(0)&1&\ddots\\ \vdots&\ddots&\ddots&\ddots\\ \end{array}\right)\,\left(\begin{array}[]{c}\psi_{1}^{[1]}\\ \psi_{2}^{[1]}\\ \psi_{3}^{[1]}\\ \vdots\end{array}\right)=$ (38)
$=-\lambda\,\left(\begin{array}[]{cccc}{\cal V}_{00}&{\cal V}_{01}&{\cal V}_{02}&\ldots\\ {\cal V}_{10}&{\cal V}_{11}&{\cal V}_{12}&\ddots\\ {\cal V}_{201}&{\cal V}_{21}&{\cal V}_{22}&\ddots\\ \vdots&\ddots&\ddots&\ddots\\ \end{array}\right)\,\left(\begin{array}[]{c}\psi_{0}(0)\\ 0\\ 0\\ \vdots\end{array}\right)\,.$ (39)
In the context of the conventional Rayleigh-Schrödinger perturbation-expansion recipe this is precisely the equation which would yield the explicit formula for coefficient $E^{[1]}$ defined in terms of the matrix elements of perturbation ${\cal V}$. Nevertheless, as long as our present unconventional unperturbed Hamiltonian is a non-diagonal (and, moreover, infinite-dimensional) matrix, we have to pay the price: The left eigenvector $\langle\!\langle\chi(0)|$ of our unperturbed Hamiltonian is not at our disposal. We cannot use it for the standard pre-multiplication of Eq (39) from the left. This means that without the knowledge of $\langle\!\langle\chi(0)|$, the first line of Eq. (39), viz., relation
$E^{[1]}={\cal V}_{00}+\psi_{1}^{[1]}/\psi_{0}(0)\,$ (40)
only enables us to extract the value of $E^{[1]}$ in the form of function of an unknown parameter $\psi_{1}^{[1]}$. This is the ambiguity which can be perceived as mimicking the unaccounted influence of the rest of the matrix elements of perturbation ${\cal V}$.
The latter formal disadvantage is partially compensated by the presence of an easily invertible triangular matrix in Eq. (39). This suggests that the role of a variable parameter could rather be played, in a partial resemblance of the EPN recipe, by the energy correction $E^{[1]}$ itself. We would then have
$\psi_{1}^{[1]}=\psi_{1}^{[1]}(E^{[1]})=(E^{[1]}-{\cal V}_{00})\psi_{0}(0)\,.$
Similarly, from the second row of Eq. (39) we would then obtain, more naturally, the value of the second wave-function component
$\psi_{2}^{[1]}=(E(0)-E_{1})\psi_{1}^{[1]}(E^{[1]})-{\cal V}_{10}\psi_{0}(0)$ (41)
etc.
7 Discussion
All in all we can conclude that in both the EPN- and IEP-related unitary-evolution scenarios the properly amended form of perturbation theory seems to be able to provide, even in its leading-order form, some explicit and useful criteria of the acceptability or unacceptability of various preselected perturbations of potential phenomenological interest.
7.1 Benign perturbations
Between the EPN and IEP alternatives there still exists a crucial difference. Indeed, in the typical EPN-related analysis our considerations usually start from our knowledge of the “physical” family of models $H(g)$. Then, the only truly difficult problem is to localize the EPN singularity, especially when the values of $N$ are not too small. In the case of the IEP singularities, in contrast, we, typically, know only the unperturbed Hamiltonian as sampled here by the IC operator. It is possible to conclude that precisely this makes the IEP-related models perceivably more difficult to study.
From a purely pragmatic point of view a source of certain optimism could really be drawn even from the above-outlined leading-order perturbation-approximation criteria. Their strength certainly lies in the possibility of identification of the “malign” IEP perturbations which would destroy the reality of the spectrum and which would make the evolution non-unitary. The complementary problem of a reliable identification of the opposite, “benign” perturbations has to be declared to stay a mathematically much more difficult open question.
Incidentally, qualitatively the same conclusions have already been obtained in the simpler EPN context. For example, in the above-mentioned study [12] of a specific Bose-Hubbard model near its EPN dynamical extreme, the authors did not insist on the reality of the spectrum and treated their mathematical results as applicable in a broader, not necessarily unitary open-system context.
In a narrower, closed-system setting, a deeper analysis and resolution of the apparent EPN-related instability paradox was described in paper [18] in which we studied the exact as well as approximate secular equations in more detail. Our ultimate conclusion was that the necessary smallness condition specifying the class of the admissible, unitarity non-violating perturbations does not involve their upper-triangular matrix part at all. In contrast, for the perturbed-EPN model in question, the lower-triangular matrix part of all of the “benign” (i.e., unitarity-compatible) perturbations has been shown to have, obligatorily, the following matrix-element-dependent form of the condition of sufficient smallness in the $\lambda\to 0$ limit,
$\lambda\,V^{(N)}_{(admissible)}=\left[\begin{array}[]{cccccc}{\lambda}^{1/2}{{\mu}}_{11}&0&\ldots&0&0&0\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr\lambda\,{\mu}_{{21}}&{\lambda}^{1/2}{{\mu}}_{22}&\ldots&0&0&0\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr{\lambda}^{3/2}\,{\mu}_{{31}}&\lambda\,{\mu}_{{32}}&\ddots&\vdots&\vdots&0\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr{\lambda}^{2}{\mu}_{{41}}&{\lambda}^{3/2}\,{\mu}_{{42}}&\ddots&\ddots&0&0\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr\vdots&\vdots&\ddots&\lambda\,{\mu}_{{{N}-1{N}-2}}&{\lambda}^{1/2}{{\mu}}_{{N}-1{N}-1}&0\\ \vskip 6.0pt plus 2.0pt minus 2.0pt\cr{\lambda}^{{N}/2}{{\mu}}_{{N}1}&{\lambda}^{({N}-1)/2}{\mu}_{{{N}2}}&\ldots&\lambda^{3/2}\,{\mu}_{{{N}{N}-2}}&\lambda\,{\mu}_{{{N}{N}-1}}&{\lambda}^{1/2}{{\mu}}_{{N}{N}}\end{array}\right]\,.$ (42)
The matrix structure (42) may be interpreted as manifesting a characteristic anisotropy and the hierarchically ordered weights of influence of the separate matrix elements because during the decrease of $\lambda\to 0$, all of the “benign” matrix-element parameters have to have bounded components $\mu_{j,k}={\cal O}(1)$. For a more explicit explanation we may rescale
$\lambda\,V^{(N)}_{(admissible)}=\lambda^{1/2}\,B(\lambda)\,V^{(reduced)}\,B^{-1}(\lambda)$ (43)
where $B(\lambda)$ would be a diagonal matrix with elements $B_{jj}(\lambda)=\lambda^{j/2}$ and where the whole reduced “benign” matrix of perturbation would be bounded, $V^{(reduced)}_{jk}={\cal O}(1)$.
[…]
7.2 IC oscillator as popular toy model
In order to elucidate the benchmark-model role of the IC IEP oscillator from a historical perspective let us recall paper [5] in which Bender with Boettcher examined a rather broad family of time-independent non-Hermitian toy-model Hamiltonians (cf. Eq. (45) in Appendix A. 2 below). They felt guided by the postulate of (antilinear) symmetry of their models,
${\cal PT}H^{(BB)}=H^{(BB)}{\cal PT}\,$ (44)
where the linear operator ${\cal P}$ can be treated as parity (causing the space reflection $x\to-x$) while ${\cal T}$ has to mimic the antilinear time reversal.
These authors proposed to treat their operators $H^{(BB)}$ as “Hamiltonians whose spectra are real and positive” so that “these ${\cal PT}-$symmetric theories may be viewed as analytic continuations of conventional theories from real to complex phase space” [5]. During the subsequent wave of development of the related mathematics it has been revealed that in the language of functional analysis the ${\cal PT}-$symmetry (44) can be re-read as the so called pseudo-Hermiticity [14] as well as a self-adjointness in the Krein space with indefinite pseudo-metric ${\cal P}$ [25, 26, 27].
A deeper mathematical insight in the class of ${\cal PT}-$symmetric models has been obtained. In the narrower context of quantum mechanics of closed systems, in contrast, the IEP-possessing IC model itself has not been assigned, up to now, any sufficiently consistent phenomenological interpretation yet [8]. Still, in retrospective, its temporary popularity was enormous. Its roots may be dated back to the Bessis’ and Zinn-Justin’s unpublished [4] but widely communicated [5] discovery that in spite of the manifest non-Hermiticity of the IC Hamiltonian its spectrum appeared to be real and bound-state-like, i.e., discrete and bounded from below.
In the extensive existing literature devoted to the study of systems with ${\cal PT}$ symmetry (cf., e.g., reviews [28, 29]), a lot of attention has been paid to the non-Hermitian but still sufficiently realistic ordinary differential Hamiltonians of the form $H=T+V$ reminiscent of the IC oscillator in which the entirely conventional kinetic-energy term $T=-d^{2}/dx^{2}$ is combined with a suitable local complex one-dimensional potential $V=V(x)$. By many authors the latter models were sampled by the field-theory-mimicking oscillator Hamiltonian (1) in which the purely imaginary form of the asymptotically growing potential is a truly puzzling mathematical curiosity.
This was also the feature which attracted a lot of attention among physicists [5, 22, 26, 28] precisely because they happened to generate the purely real, discrete and non-negative (i.e., hypothetically, observable and bound-state-like) energy spectra. Still, the ultimate verdict by mathematicians [1, 8] was discouraging because they proved that the IC Hamiltonian cannot be assigned any isospectral self-adjoint avatar $\mathfrak{h}({{t}})$ or acceptable physical inner-product metric [1]. Thus, the rigorous mathematical analysis finally led to the loss of some of the most optimistic phenomenological expectations.
8 Summary
Not too surprisingly, the highly desirable proofs of the so called unbroken form of ${\cal PT}-$symmetry (in which, by definition [22], the spectrum remains real) appeared to be, in numerous applications, strongly model-dependent. There seemed to be no universal criteria guaranteeing the existence of the unbroken ${\cal PT}-$symmetry in dependence on a suitable measure or degree of the non-Hermiticity of the Hamiltonian.
During the preparation of our present study we came to the conclusion that the lack of a deeper understanding of correspondence between the (apparent) non-Hermiticity and (hidden) unitarity might have been caused by an overambitious generality of the choice of the models in the literature. For this reason we decided to narrow the scope of our analysis and we decided to restrict our attention just to the extremely non-Hermitian Hamiltonians which would lie very close to their EPN or IEP singularity.
In our present paper we explained that and how such a decision enabled us not only to pick up a rather natural measure of the non-Hermiticity (characterized simply by the inverse distance of $g$ from its exceptional value) but also to formulate a well-defined project in which we developed and applied, consequently, some innovative and suitable perturbation-approximation techniques. In its framework we managed to show that the unbroken ${\cal PT}-$symmetry of our models can really survive inside an open parametric domain, on a point of boundary of which our measure of non-Hermiticity reaches its maximum. The latter point has been shown to coincide either with the Kato’s [7] exceptional point of a finite order $N$ or with its hypothetical IEP analogue.
Such an approach has been found productive. Using certain slightly modified techniques of perturbation theory of linear operators we found that, paradoxically, the restriction of attention to the smallest vicinity of the singularity (in which the Hamiltonians become maximally non-Hermitian) leads to a remarkable simplification of the perturbation-approximation constructions. The special, “exceptional” non-diagonalizable operators appeared to be eligible as unperturbed Hamiltonians in spite of their being singular and unacceptable as observables. In their vicinity (denoted as ${\cal D}$) their diagonalizability as well as the observability status were re-established.
Constructively we managed to clarify also some of the consequences. The emergence of qualitative as well as quantitative EPN - IEP parallels helped us to complement and understand better the twelve years old disproof of the internal mathematical consistency of the IC IEP quantum oscillator [1]. Such a clarification can be perceived as being of a fundamental importance in quantum theory. Indeed, potentially, most of our observations might immediately be extended also to many other currently popular non-Hermitian quantum models.
Appendices
A. 1. Paradox of stable bound states in complex potentials
For a long time it was believed that the locality of the real and confining potential is so strongly restrictive a constraint that the loss of the reality of $V(x)$ (i.e., of the self-adjointness of the Hamiltonian in any standard Hilbert space of states, i.e., say, in $L^{2}(\mathbb{R})$) would immediately imply the loss of the reality of the spectrum, i.e., the loss of the observability of the quantum system in question.
In 1998, in the Bender’s and Boettcher’s pioneering letter [5] the latter belief has been strongly opposed. Using a combination of methods these authors argued that also the spectrum generated by multiple manifestly complex  local interaction potentials $V(x)$ still appears to be strictly real and discrete, i.e., fully compatible with the conventional postulates of quantum mechanics of the stable and unitary bound-state quantum systems.
Subsequently, the proposed amendment of the model-building paradigm has widely been accepted. For various complex $V(x)$s, rigorous [6, 30] as well as numerical [31] proofs of the reality of the spectra were found and attributed to a certain “hidden form of Hermiticity” of the underlying Hamiltonians (cf., e.g., a few earlier review papers [14, 22] for details).
A return to older literature (cf., e.g., review [10]) revealed that the compatibility of the unitarity of the evolution with a manifest non-Hermiticity of the interaction can be given a comparatively elementary explanation because whenever the Hamiltonian $H$ in question has a real and discrete spectrum, it may be safely self-adjoint  with respect to another, “correct”, ad hoc  inner product, i.e., in a modified, “physical” Hilbert space ${\cal H}_{phys}$. Simultaneously, it may make sense to stay working in the initial and more user-friendly Hilbert space ${\cal H}_{math}$ which remains “unphysical” (i.e., formally non-equivalent) but, for some reasons, preferred.
Many years ago many people really studied various toy models of such a type, characterized by the interaction which appeared manifestly non-Hermitian  with respect to a conventional inner product. The scope of such – mostly, numerical – attempts ranged from very pragmatic Dyson-inspired analyses of non-relativistic many-body systems [32, 33] up to the abstract, methodically motivated considerations concerning the applicability of non-Hermitian models in the relativistic quantum field theory [4].
In the latter context, the Bender’s and Boettcher’s results [5] proved particularly inspiring and made the idea popular. In parallel, the most elementary IC model appeared to represent a challenge in mathematics, leading, i.a., to a rigorous proof of the reality of its spectrum by Dorey et al [6]. For this reason the model served, for many years, as a benchmark methodical guide which inspired several new developments in relativistic quantum field theory [34] as well as in multiple other phenomenologically oriented subdomains of modern physics [29, 35].
Last but not least, Bender with Boettcher extended the spectrum-reality conjecture to a broad class of potentials $V^{(BB)}(x)=({\rm i}x)^{\delta}\,x^{2}$ with arbitrary non-negative $\delta\in(0,\infty)$ and with $x$ lying on a complex contour [5]. All of these results caused the growth of the popularity of the innovative reformulation of quantum physics of unitary systems admitting manifestly non-Hermitian Hamiltonians among physicists. This, not too surprisingly, appeared paralleled by a criticism by mathematicians who referred, e.g., to the existence of counterexamples with pathological properties [2]. Incidentally, some of these counterexamples were even already known, many years earlier, to Dieudonné [9]).
Fortunately, a temporary resolution of the conflict has been found in a rediscovery and return to a half-forgotten but still fully relevant older review paper by Scholtz et al [10]. In it, most of the objections by mathematicians were circumvented by an ad hoc  technical assumption that one is only allowed to consider the non-Hermitian Hamiltonians (as well as any other candidates for observables) which are, as operators in Hilbert space, bounded: cf. also several mathematically oriented reviews in [26] in this respect. Thus, one may call such a mathematically consistent version of the theory quasi-Hermitian quantum mechanics.
A. 2. Beyond the imaginary cubic-oscillator potential
The above-mentioned Bender’s and Boettcher’s choice of the illustrative stationary non-Hermitian one-dimensional (i.e., ordinary differential) Hamiltonians
$H^{(BB)}=-\frac{d^{2}}{dx^{2}}+V_{(\delta)}(x)\,,\ \ \ \ V_{(\delta)}(x)=({\rm i}x)^{\delta}\,x^{2}\,,\ \ \ \ \delta\in(0,\infty)\,$ (45)
has been motivated by their conjecture that the ubiquitous requirement of the self-adjointness of the observables might be criticised as over-restrictive. They proposed that one should consider a broader class of Hamiltonians $H$ for which the conventional condition of self-adjointness becomes replaced by the property called ${\cal PT}-$symmetry (cf. Eq. (44)) alias  ${\cal P}-$pseudo-Hermiticity,
${\cal P}H^{(BB)}=\left[H^{(BB)}\right]^{\dagger}{\cal P}\,.$ (46)
Under the latter assumption (cf. also the comments in [25, 36]) Bender with Boettcher assumed that the role of the guarantee of the reality of the spectrum of the bound-state energies (i.e., in principle, of their observability) can be relegated from the conventional Hermiticity to the ${\cal PT}-$symmetry of the system whenever such a symmetry remains spontaneously unbroken [22].
In applications the choice of ${\cal PT}-$symmetric Hamiltonians appeared strongly influenced by a tacit reference to the principle of correspondence due to which $H$ is assumed split into its kinetic-energy component $H_{kin}$ and a suitable interaction term $H_{int}$. Moreover, the analysis is often restricted just to the single-particle one-dimensional motion with conventional $H_{kin}\sim-d^{2}/dx^{2}$ and with a suitable local-interaction form of $H_{int}\sim V(x)$.
This choice has already been recommended by Bender with multiple collaborators (cf. review [22]). They emphasized that the study of various non-Hermitian but ${\cal PT}-$symmetric quantum models with real spectra can be perceived as motivated by quantum field theory. In this context a key role is played, in a way proposed by Bessis with collaborators [4], by the imaginary cubic (IC) potential $V^{(IC)}(x)={\rm i}x^{3}$. For this reason, Siegl with Krejčiřík [1] turned their attention to the IC Hamiltonian (1), having revealed that such a model suffers of unpleasant pathologies.
These pathologies appeared only too serious to be ignored. After all, Siegl with Krejčiřík only rediscovered the Dieudoné’s older claim that for such a Hamiltonian “there is for instance no hope of building functional calculus that would follow more or less the same pattern as the functional calculus of self-adjoint operators” [9]. Siegl with Krejčiřík also listed several “pathological properties of non-self-adjoint $H^{(IC)}$” and they offered a rigorous proof that these features of the IC model find a close formal analogue in the Kato’s EPNs.
Thus, the popular toy-model operator $H^{(IC)}$ became disqualified as a candidate for quantum Hamiltonian (see also an independent reconfirmation of the skepticism, say, in [12, 13] and, after all, also in the very first line of the abstract of the comprehensive review [14] requiring the diagonalizability of the observables). Still, several attempts were made to replace $H^{(IC)}$ by a suitable regularized alternative. Typically, the regularization has been sought in a truncation of the real line of $x$ (cf. [37]). Unfortunately, one can hardly speak about a successful resolution of the problem because one form of unacceptability was merely replaced by another one, viz., by the complexification of the spectrum.
A. 3. Beyond the stationary quasi-Hermitian models
What is most characteristic for the applications of quantum mechanics in the so called Schrödinger picture [15] is the observability of the generator of the evolution of wave functions called “Hamiltonian”. In most applications it is required self-adjoint in the preselected Hilbert space ${\cal H}$. Then, according to Bender and Boettcher [5], the robust nature of the reality of its eigenvalues (representing, in many models, just the discrete bound-state energies) can be perceived as a weakness of the approach. Indeed, once we prepare, at an initial time $t=0$, the system in a pure state (alias  “phase”) described by a ket-vector $|\psi(0)\rangle\in{\cal H}$, we discover that the “phase” (defined by the specific set of observable aspects) cannot be changed by the evolution.
This feature would make the description of quantum phase transitions impossible. Fortunately, a change of the “phase” (i.e., e.g., an abrupt loss of the observability of the time-dependent bound-state energies) has recently been rendered possible after a conceptually straightforward transition from the Schrödinger-picture (SP) approach to a formally equivalent, albeit technically more complicated non-Hermitian interaction picture (NIP, [38]).
Some of the consequences of such a change of paradigm become relevant also for an appropriate understanding and treatment of the IEP-related considerations. Due to the lack of space for an exhaustive analysis of this problem, let us only briefly mention that in the traditional and most popular SP framework of conventional textbooks the unitary evolution of a closed quantum system is just being described in a unique preselected Hilbert space ${\cal H}$. People also often accept multiple additional ad hoc  simplifying assumptions, with the most popular one concerning the above-mentioned generator of evolution of wave functions (say, $G_{(textbook)}(t)=\mathfrak{h}(t)$) and requiring its self-adjointness in ${\cal H}$,
$\mathfrak{h}(t)=\mathfrak{h}^{\dagger}(t)\,.$ (47)
Another such a traditional simplification concerns the inner product in ${\cal H}$ which is assumed time-independent [10, 14].
In the generalized NIP framework, in contrast, one has to consider Schrödinger equation
${\rm i}\frac{d}{dt}\,|\psi(t)\rangle=G(t)\,|\psi(t)\rangle\,,\ \ \ \ \ |\psi(t)\rangle\in{\cal H}\,$ (48)
in which the generator $G(t)$ need not represent an observable [39, 40, 41, 42, 43]. From the purely phenomenological point of view such a generalization is useful. In a way reflecting the widespread knowledge of the above-sampled differential-operator benchmark models (45) it is still possible to introduce the energy-representing observables
$H(t)=\triangle+V(x,t)$ (49)
which are not only non-Hermitian and manifestly time-dependent but also different from the generator $G(t)$. In this setting the freedom of choice between the SP or NIP framework only means that one treats the time-dependence of our observables as inessential or essential, respectively.
In practice we usually insist on the standard phenomenological and probabilistic interpretation and, in particular, on the observable energy status of the specific operator (49). Thus, we have to keep in mind that at least some of the most popular differential operators cannot be used as benchmark models without hesitation. For this reason, a consequent constructive realization of description of the phenomenon of a genuine quantum phase transition remains to be a task for the future development of the theory.
A. 4. The question of the unitary-evolution accessibility of EPNs
Due to the degeneracy of the unperturbed energy spectrum in EPN limit the $N-$plet of the perturbed-energy roots of the corresponding secular equation (cf., e.g., the bound-state energy roots $\epsilon_{n}=\epsilon_{n}(\lambda)$ of Eq. (20) with $n=1,2,\ldots,N$) need not necessarily be all real and, hence, representing observable quantities. In Refs. [12, 17], for example, even some of the approximate leading-order roots were found complex. This observation can be reinterpreted as indicating that even in an immediate EPN vicinity even the bounded perturbations may still be reclassified, in unitary theory, as “inadmissibly large”, forcing the system to perform an abrupt quantum phase transition.
Within quantum mechanics of unitary, closed systems in its quasi-Hermitian formulation a key to the suppression of such a quantum catastrophe lies in the construction of a correct physical inner product in Hilbert space [10]. Still, many of the truly realistic applications of the quasi-Hermitian operators may remain, in the model-building context, counterintuitive. In particular, the doubts emerge in virtually all of the tentative quasi-Hermitian descriptions of the phenomenon of quantum phase transition because, traditionally, all of such processes have been treated as non-unitary, requiring an ad hoc  effective-operator approach [44].
A feasible way out of the apparent deadlock is offered by the quasi-Hermitian quantum models in which a given observable with real spectrum (say, $\Lambda(t)$) is non-Hermitian. In such a theory (cf., e.g., its reviews [10, 14, 20, 22, 26]) the condition of self-adjointness of $\Lambda(t)$s survives “in disguise”, being replaced by a formally equivalent quasi-Hermiticity condition in another Hilbert space,
$\Lambda^{\dagger}(t)\,\Theta(t)=\Theta(t)\,\Lambda(t)\,.$ (50)
The assumption of the time-dependence of the related inner-product-metric $\Theta(t)$ opens then the possibility of reaching a singularity via unitary evolution.
In such a case the collapse is simply rendered possible by the coherent, simultaneous loss of the existence of the time-dependent inter-twiner $\Theta(t)$ in the critical limit of $t\to t^{(EPN)}$ or $t\to t^{(IEP)}$. One could also say that the realization of the whole process of the change of phase, i.e., of the loss of the observability of some of the measurable characteristics (i.e., of the loss of the quasi-Hermiticity of $\Lambda(t)$) is to be mediated by the metric $\Theta(t)$ in (50) which becomes, in the limit, non-invertible and, in fact, just a rank-one operator [11].
The emergence of a fully explicit conflict between the constructive feasibility and mathematical consistency can be traced back to the year 2012 and paper [1] in which Siegl with Krejčiřík disproved the acceptability of a broad class of the currently popular non-Hermitian but observable Hamiltonians. The impact of the disproof was truly destructive. The currently widespread belief in the benchmark role of many non-Hermitian but still observable Hamiltonians with real spectra has been shattered.
In parallel, the doubts were also thrown upon the acceptability of the specific benchmark ordinary-differential (i.e., one-dimensional and mathematically still sufficiently user-friendly) non-Hermitian candidates for the energy-representing Hamiltonians decomposed into their two intuitively plausible (and stationary as well as non-stationary) kinetic- plus potential-energy components,
$H=-\frac{\hbar}{2m(x)}\frac{d^{2}}{dx^{2}}+V(x)\neq H^{\dagger}\,.$ (51)
The no-go theorems of paper [1] (see also [3] for further details) seemed to return us back to the older methodical analyses in which the deepest source of the mathematical difficulties has been attributed to the unbounded-operator nature of the most popular differential operators as sampled by Eq. (51) (cf., e.g., [9, 10, 45]).
In applications, paradoxically, the latter disproofs and skepticism motivated a rapid increase of interest in the study of the so called open quantum systems [35]. In such a very traditional context an enormous acceleration of the progress (say, in an innovated understanding of the dynamics of resonances) has been achieved due to the successful applications of the new methods of the solution of the non-Hermitian versions of the Schrödinger-like evolution equations. At present, multiple branches of physics were enriched by these tendencies, including even the non-quantum ones [28, 29].
A. 5. A note on the broader quantum-physics framework
Even in the context of quantum physics, paradoxically, the intensification of interest in non-Hermitian Hamiltonians of closed systems accelerated the progress in the development of the dedicated mathematical methods and, in particular, in the understanding of non-Hermitian operators with the spectrum which was not real. Paradoxically, these developments redirected attention of a part of physics community back to the traditional models in which the meaningful spectra were allowed to be complex.
Our lasting interest in unitary quantum mechanics using hiddenly Hermitian observables is partly motivated, in a way documented in paper [46], by the possibility of mimicking the processes of quantum phase transitions. We believe that such a direction of analysis must necessary profit from the very recently introduced combinations of the requirements of the manifest non-Hermiticity (and, especially, of its hiddenly Hermitian forms called quasi-Hermiticity [9, 10]) with some other suitable model-building options and auxiliary technical assumptions like the time-dependence of the operators and/or their ${\cal PT}-$symmetry (cf. [5, 22, 36]) or factorization (cf., e.g., [36, 47, 48]).
The main stream of our considerations remained restricted to the context of quantum physics and quasi-Hermitian dynamics in which we insisted on the compatibility of our models with all of the basic principles of quantum mechanics of the so called closed and unitary systems admitting the standard probabilistic interpretation. This does not mean that the scope of the theory and of its applications cannot be much broader, in principle at least. Innovations may be obtained in the representation of the states as well as of their observable characteristics.
In the spirit of multiple relevant recent reviews this goal can be achieved by the various physical reinterpretations of the parameter-dependent Hamiltonians $H(g)$. Even when we only admit, in Schrödinger picture, its observable-energy interpretation, it is still worth returning to the Dyson’s treatment [32] of such an operator as the one which is isospectral with its self-adjoint avatar $\mathfrak{h}({{g}})$,
$H({{g}})\to\mathfrak{h}({{g}})=\Omega({{g}})\,H({{g}})\,\Omega^{-1}({{g}})=\mathfrak{h}^{\dagger}({{g}})\,.$ (52)
In this manner, even the metric $\Theta$ itself acquires an entirely new meaning of the mere product
$\Theta(g)=\Omega^{\dagger}{({g})}\,\Omega({{g}})$ (53)
of the so called Dyson’s maps which are non-unitary and related to the conventional quantum physics avatar $\mathfrak{h}({{g}})$ rather than to the non-Hermitian upper-case Hamiltonian itself.
In this spirit, Dyson introduced and treated the mappings $\Omega$ in (52) as certain variationally motivated ad hoc  multiparticle correlations. In contrast, Buslaev with Grecchi [36] offered and formulated another point of view by which these operators represent just an isospectrality-equivalence transition from a Hilbert space which is unphysical to another Hilbert space which is physical. During such a transition it is possible to distinguish and cover both the open quantum systems (in which one describes resonances) and the closed quantum systems (in this case, in loc. cit., Buslaev with Grecchi paid their attention to the quartic anharmonic oscillators).
In the latter, newer and less traditional case one has to construct the auxiliary metric as product (53). This just amends the inner product in the mathematically friendlier and computationally preferred but manifestly unphysical Hilbert space ${\cal K}$,
$\langle\psi_{1}|\psi_{2}\rangle_{in\ physical\ space}=\langle\psi_{1}|\Theta\,|\psi_{2}\rangle_{in\ mathematical\ space}\,.$ (54)
In this notation the obligatory condition of the Hermiticity of $\mathfrak{h}({{g}})$ (cf. Eq. (52)) becomes translated into the equally obligatory condition of the quasi-Hermiticity of $H({{g}})$ in the mathematical Hilbert space,
$H^{\dagger}({{g}})\,\Theta({{g}})=\Theta({{g}})\,H({{g}})\,.$ (55)
Thus, for a preselected Hamiltonian $H(g)$ with real spectrum, its acceptability as a closed-system observable will be guaranteed either by the Hermiticity of its $\Omega-$transformed isospectral avatar or, equivalently, by the $\Theta-$quasi-Hermiticity property of $H(g)$ itself.
A. 6. Final note on the notation and outlook
In the literature devoted to the models using quasi-Hermitian observables the notation conventions did not unite yet. Thus, the Hilbert-space metric (which we decided to denote by the upper-case Greek symbol $\Theta$) can be found denoted as $T$ (which does not mean time reversal, cf. equation Nr. (2.2) in one of the oldest reviews [10]) or as subscripted lower-case Greek $\eta_{+}$ (cf. equation Nr. (52) in one of the more modern reviews [14]) or as $\exp(-Q)$ (cf. the 2006 paper [49]) or as $\rho$ (cf. [50]) or by the letter $G$ (cf. Tables Nr. I and II in [51]), etc.
The notion of EPs (exceptional points) of our present interest emerged within the strictly mathematical theory of linear operators [7]. It played there a key role in the rigorous analysis of the criteria of convergence of perturbation series. In the context of physics, the notion was less well known, being called there the Bender-Wu singularity [52] etc. Independently, this notion has only been found important for physicists [53, 54], especially during the last twenty years, viz., during the growth of interest in the role played by the non-self-adjoint operators in several (i.e., not always just quantum) branches of phenomenology [28, 29].
During the early stages of development of the latter innovative approach the role of a benchmark illustrative example has been played by the IC (imaginary cubic) differential-operator Hamiltonian of Eq. (1). Later, as we already explained above, such a choice of illustration proved to be a bit unfortunate. For proof we cited Siegl and Krejčiřík, [1] who emphasized that, in the formal sense, the obstacles imposed by the loss of the Riesz-basis diagonalizability of the IC Hamiltonian are “much stronger than” those imposed by ”any exceptional point associated with finite Jordan blocks”.
The related concept of asymptotic IEP was not only very new but also rather elusive. Even its definition as provided by the authors was just implicit (see sections IID, IIIC and IV in [1]). An explanation is that their message was aimed, first of all, at the community of physicists for which the IEP IC oscillator model served as a heuristic “fons and origo” of what has been widely accepted as ${\cal PT}-$symmetric quantum mechanics. The same authors also emphasized that the properties of $H^{(IC)}$ “are essentially different with respect to self-adjoint Hamiltonians, for instance, due to spectral instabilities”. Thus, the main IEP-related result of [1] (viz., the proof of the existence of an IEP anomaly in the IC model) was finally formulated as an observation that “there is no quantum-mechanical Hamiltonian associated with it via …similarity transformation”.
The latter conclusion was revolutionary and opened a number of new questions concerning the necessity of finding “new directions in physical interpretation” of the model. In our present paper we, perhaps, threw new light on the issue, with a rather sceptical conclusion that the currently unresolved status of the twelve years old conceptual task of the interpretation of the IEP-related instabilities does not seem to have an easy resolution, indeed.